# CMSO-transducing tree-like graph decompositions

### v2

@@ 3.1 Inner node representatives
Proof 3.10.
We proceed by contradiction and thus assume $S\neq S^{\prime}$. Let $s\in S\setminus S^{\prime}$. By Lemma 3.7, the number of children $c$ of $s$ for which the set $(A\cup B)\cap V(T_{c})$ has odd size is $2$ since $s\in S$, while it should also be less than $2$ since $s\notin S^{\prime}$. A contradiction.
@@
Figure 2: Illustration of a bi-colouring $(A,B)$ identifying some set $S\subseteq V\setminus L$ in a binary tree $T$. Leaves from $A$ are coloured blue, leaves from $B$ are coloured red, and inner nodes from $S$ are marked purple. Furthermore, the two paths connecting an inner node $s\in S$ to its $A$- and $B$-representatives are coloured blue and red, respectively.
@@
Proof 3.12.
If $T_{s}$ has height 00, then $s$ is a leaf and taking $t=s$ trivially gives the expected path. Otherwise, $s$ is an inner node and, because $X$ is thin, $s$ has at least one child $c_{s}$ not belonging to $X$. By induction, there is a path from some leaf $t\in V(T_{c_{s}})$ to $c_{s}$ avoiding $X$ and, since $s\notin X$, this path could be extended into a path from $t$ to $s$ avoiding $X$.
The previous lemma allows to identify every thin set.
Lemma 3.13.
@@
Proof 3.14.
We proceed by induction on the size of $X$. If $X=\emptyset$, the result is trivial. Let $n\in\mathbb{N}$ and suppose that for every thin set of size $n$ there exists a pair of injections identifying it with unique request. Let $X$ be a thin set of size $n+1$, and let $s\in X$ be of minimal depth. Clearly, $X\setminus\{s\}$ is thin and thus there exists, by induction, a pair $(\pi,\sigma)$ of injections from $X\setminus\{s\}$ to $L$ identifying $X\setminus\{s\}$ with unique request. Since $X$ is thin and $s\in X$, we can find two distinct children $c_{a}$ and $c_{b}$ of $s$ not belonging to $X$ [^5] Remember that every inner node of $T$ has at least two children (including possible leaves). Then, by Lemma 3.11, there exists a leaf $a\in V(T_{c_{a}})$ (resp. a leaf $b\in V(T_{c_{b}})$) such that the path from $a$ to $c_{a}$ (resp. from $b$ to $c_{b}$) avoids $X$. In particular, for each node $y$ along these paths, since $y$ has no ancestor that belongs to $X$ but $s$, $y$ is not requested in $(\pi,\sigma)$. Hence, extending $\pi$ (resp. $\sigma$) in such a way that, besides mapping each $x\in X\setminus\{s\}$ to $\pi(x)$ (resp. to $\sigma(x)$), it maps $s$ to $a$ (resp. to $b$), we obtain a pair $(\hat{\pi},\hat{\sigma})$ of injections from $X$ to $L$ that identifies $X$ with unique request.
A family $F=(A_{1},B_{1}),\ldots,(A_{n},B_{n})$ of bi-colourings identifies a set $S\subseteq V\setminus L$, if there exists a partition $(S_{1},\ldots,S_{n})$ of $S$ such that, for each $i\in[n]$, $(A_{i},B_{i})$ identifies $S_{i}$. Whenever $S=V\setminus L$ we say that $F$ identifies $T$. A collection of subsets of $V\setminus L$ is thin if each of its subsets is thin. We now show that there exists a thin $4$-partition.
Lemma 3.15.
@@
Proof 3.18.
The result immediately follows from Lemmas 3.13 and 3.15.
Figure 3: Illustration of a partition of the inner nodes into 4 thin sets indicated by different shapes/fillings. The four colourings identifying each thin set are indicated in the table below the leaves.
3.2 The transduction
The goal of this section is to prove Theorem 3.1, that is, to design a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction that produces the laminar tree induced by an input laminar set system. We fix a laminar set system $(U,\mathcal{S})$, represented by a $\{\mathsf{SET}\}$-structure $\mathbb{S}$. Before proving the theorem, we make the following basic observation. On $\mathbb{S}$, we can define two $\operatorname{MSO}$-formulas $\mathsf{desc}(X,Y)$ and $\mathsf{child}(X,Y)$ expressing that in the laminar tree induced by $\mathcal{S}$, $X$ and $Y$ are nodes and $X$ is a descendant or a child of $Y$, respectively:
@@ 3.2 The transduction
Lemma 3.19.
[…]
Proof 3.20.
First, we define a formula $\phi_{A,B}(X)$ that, under the above assumption, is satisfied exactly when $X$ belongs to $S$. According to Lemma 3.7, this happens if and only if $X$ is a node of $T$ (i.e.$\mathsf{SET}(X)$ is satisfied) and there exists $a\in X\cap A$ and $b\in X\cap B$ such that for each child $Z$ of $X$, $\{a,b\}\nsubseteq Z$ and the set $(Z\setminus\{a,b\})\cap(A\cup B)$ has even size. This property is easily expressed in $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$, using the $\operatorname{MSO}$-formula $\mathsf{child}(X,Y)$ defined previously, as well as the predicate $\mathsf{SET}$:
@@
[…]
@@
Proof 3.21 (Proof of Theorem 3.1).
[…]
  1. 1.
    Guess a family of four bi-colourings $(A_{i},B_{i})_{i\in[4]}$ identifying $T$ (which exists by Corollary 3.17).
  2. 2.
@@
  3. 3.
[…]
  4. 4.
[…]
4 Transducing modular decompositions
@@
4.1 Transducing weakly-partitive trees
@@
A member of a set system $(U,\mathcal{S})$ is said to be strong if it does not overlap any other set from $\mathcal{S}$. The sub-family $\mathcal{S}_{!}$ of strong sets of $\mathcal{S}$ is thus laminar by definition. Hence, it induces a laminar tree $T$. By extension, we say that $T$ is induced by the set system $(U,\mathcal{S})$ (or simply by $\mathcal{S}$). The next result extends Theorem 3.1, by showing that $T$ can be $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduced from $\mathcal{S}$.
Lemma 4.1.
Let $\Sigma$ be an extended vocabulary, including a set unary predicate name $\mathsf{SET}$ and not including the binary relational symbol $\mathsf{ancestor}$. There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each set system $(U,\mathcal{S})$ represented as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ and inducing the laminar tree $T$ with ${L(T)=U}$, and for each $\Sigma$-structure $\mathbb{A}$ with $\mathbb{S}\sqsubseteq\mathbb{A}$, $\tau(\mathbb{A})$ is non-empty and every output in $\tau(\mathbb{A})$ is equal to some $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ representing $T$.
Proof 4.2.
@@
Hence, we can design an $\operatorname{MSO}$-interpretation that outputs the $\Sigma\cup\{\mathsf{SET!}\}$-structure corresponding to $\mathbb{A}$ equipped with the set unary predicate $\mathsf{SET!}$ that selects strong members of $\mathcal{S}$. Thus, up to renaming the predicates $\mathsf{SET!}$ and $\mathsf{SET}$, by Theorem 3.1, we can produce, through a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction, the $\Sigma\cup\{\mathsf{ancestor}\}$-structure $\mathbb{A}\sqcup\mathbb{T}$ where $\mathbb{T}$ is the tree-structure modeling the laminar tree $T$ induced by $\mathcal{S}_{!}$ with $L(T)=U$ (once the output obtained, the interpretation drops the set predicate $\mathsf{SET!}$ which is no longer needed).
Clearly, the laminar tree $T$ of a weakly-partitive set system $(U,\mathcal{S})$ does not characterize $(U,\mathcal{S})$. However, as shown by the below theorem, a labeling of its inner nodes and a controlled partial ordering of its nodes are sufficient to characterize all the sets of $\mathcal{S}$. For $Z$ a set equipped with a partial order $<$ and $X$ a subset of $Z$, we say that $X$ is a $<$-interval whenever $<$ defines a total order on $X$ and for every $a,b\in X$ and every $c\in Z$, $a<c<b$ implies $c\in X$.
Theorem 4.3 ([7, 25]).
@@
Let $\mathcal{S}$ be a weakly-partitive family, $\mathcal{S}_{!}$ be its subfamily of strong sets, and $T$ be the laminar tree it induces. There exists a total labeling function $\lambda$ from the set $V(T)\setminus L(T)$ of inner nodes of $T$ to the set $\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$, and, for each inner node $t\in\lambda^{-1}(\mathsf{linear})$, a linear ordering $<_{t}$ of its children, such that every inner node having exactly two children is labeled by $\mathsf{degenerate}$ and the following two conditions are satisfied:
@@
  • for each $X\in\mathcal{S}\setminus\mathcal{S}_{!}$, there exists $t\in V(T)$ and a subset $\mathcal{C}$ of children of $t$ such that $X=\bigcup_{c\in\mathcal{C}}L(T_{c})$ and either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is a $<_{t}$-interval, or $\lambda(t)=\mathsf{degenerate}$;
@@
  • conversely, for each inner node $t$ and each non-empty subset $\mathcal{C}$ of children of $t$, if either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is a $<_{t}$-interval, or $\lambda(t)=\mathsf{degenerate}$, then $\bigcup_{c\in\mathcal{C}}L(T_{c})\in\mathcal{S}$.
@@
Furthermore, $T$ and $\lambda$ are uniquely determined from $\mathcal{S}$, and, for each inner node $t$ of $T$ labeled by $\mathsf{linear}$, only two orders $<_{t}$ are possible, one being the inverse of the other (indeed, inverting an order $<$ does preserve the property of being a $<$-interval). Hence, every weakly-partitive family $\mathcal{S}$ is characterized by a labeled and partially-ordered tree $(T,\lambda,<)$ where $T$ is the laminar tree induced by the subfamily $\mathcal{S}_{!}$ of strong sets of $\mathcal{S}$, $\lambda:V(T)\setminus L(T)\to\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$ is the labeling function, and $<$ is the partial order $\bigcup_{t\in\lambda^{-1}(\mathsf{linear})}<_{t}$ over $V(T)$. As, up-to inverting some of the $<_{t}$ orders, $(T,\lambda,<)$ is unique, we abusively call it the weakly-partitive tree induced by $\mathcal{S}$. Conversely, a weakly-partitive tree characterizes the unique weakly-partitive set system which induced it.
@@
We naturally model a weakly-partitive tree $(T,\lambda,<)$ of a partitive set system $(U,\mathcal{S})$ by the $\{\mathsf{ancestor},\mathsf{degenerate},\mathsf{betweeness}\}$-structure $\mathbb{T}$ of universe $U_{\mathbb{T}}=V(T)$ such that $\langle V(T),\mathsf{ancestor}_{\mathbb{T}}\rangle\sqsubseteq\mathbb{T}$ models $T$ with ${L(T)=U}$, $\mathsf{degenerate}_{\mathbb{T}}$ is a unary relation which selects the inner nodes of $T$ of label $\mathsf{degenerate}$, i.e., ${\mathsf{degenerate}_{\mathbb{T}}=\lambda^{-1}(\mathsf{degenerate})}$, and $\mathsf{betweeness}_{\mathbb{T}}$ is a ternary relation selecting triples $(x,y,z)$ satisfying $x<y<z$ or $z<y<x$. (Although it is possible, through a non-deterministic $\operatorname{MSO}$-transduction, to define $<$ from $\mathsf{betweeness}_{\mathbb{T}}$, the use of $\mathsf{betweeness}_{\mathbb{T}}$ rather than $<_{\mathbb{T}}$ ensures unicity of the output weakly-partitive tree.) The inner nodes of $T$ that are labeled by $\mathsf{linear}$ can be recovered through an $\operatorname{MSO}$-formula as those inner nodes whose children are related by $\mathsf{betweeness}$. The inner nodes labeled by $\mathsf{prime}$ can be recovered through an $\operatorname{MSO}$-formula as those inner nodes that are labeled neither by $\mathsf{degenerate}$ nor by $\mathsf{linear}$. Using Theorem 4.3, it is routine to design an $\operatorname{MSO}$-transduction which takes as input a weakly-partitive tree and outputs the weakly-partitive set system which induced it. The inverse $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction is the purpose of the next result.
Theorem 4.4.
@@
Proof 4.5.
By Lemma 4.1, we have a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which on inputs $\mathbb{S}$ outputs the $\{\mathsf{SET},\mathsf{ancestor}\}$-structure $\mathbb{S}\sqcup\mathbb{T^{\prime}}$, where $\mathbb{T^{\prime}}$ is the $\{\mathsf{ancestor}\}$-structure modeling $T$, with $L(T)=U=U_{\mathbb{S}}$. It thus remains to define a second $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which outputs $\mathbb{T}$ from $\mathbb{S}\sqcup\mathbb{T^{\prime}}$. Actually, an $\operatorname{MSO}$-interpretation is enough. Indeed, we can define an $\operatorname{MSO}$-formula $\phi_{\mathsf{degenerate}}$ which selects the inner nodes of $T$ that are labeled $\mathsf{degenerate}$ by $\lambda$, and an $\operatorname{MSO}$-formula $\phi_{\mathsf{betweeness}}$ which selects triples $(x,y,z)$ of elements such that $x<y<z$ or $z<y<x$. In order to define both formulas, we use a functional symbol $\mathsf{leafset}(t)$ which is interpreted as the set $L(T_{t})$, and is clearly $\operatorname{MSO}$-definable on $\{\mathsf{ancestor}\}$-structures such as $\mathbb{T}$. Indeed, an element $z$ belongs to $\mathsf{leafset}(t)$ if and only if it is a leaf and has $t$ as ancestor.
According to Theorem 4.3, an inner node $t$ is labeled $\mathsf{degenerate}$ if and only if, for every two children $s_{1},s_{2}$ of $t$ the set $L(T_{s_{1}})\cup L(T_{s_{2}})$ is in $\mathcal{S}$. We can define this in $\operatorname{MSO}$ using the following formula:
@@
Now, according to Theorem 4.3, we have $x<y<z$ or $z<y<x$ exactly when $x$, $y$, and $z$ are three distinct children of some node $t$ and the three following properties are satisfied:
  1. 1.

### v4

@@ 3.1. Inner node representatives
Proof 3.4.
We proceed by contradiction and thus assume $S\neq S^{\prime}$. Let $s\in S\setminus S^{\prime}$. Since $s\in S$ by 7(3), there are two leaves $a\in A\cap V(T_{c})$ and $b\in B\cap V(T_{c})$ such that, for each child $c$ of $x$, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|(B\setminus\{b\}\cap V(T_{c}))\right|$ and $\{a,b\}\nsubseteq V(T_{c})$. Let $c_{a}$ and $c_{b}$ be the two children of $s$ with $a\in V(T_{c_{a}})$ and $b\in V(T_{c_{b}})$. Since $\{a,b\}\nsubseteq V(T_{c})$ for every child $c$ of $s$, we get $c_{a}\neq c_{b}$. In particular, $\left|A\cap V(T_{c_{a}})\right|=\left|(B\cap V(T_{c_{a}}))\right|+1$, $\left|A\cap V(T_{c_{b}})\right|+1=\left|(B\cap V(T_{c_{b}}))\right|$ and $\left|A\cap V(T_{c})\right|=\left|(B\cap V(T_{c}))\right|$ for every other children $c$ of $s$. Since $A$ and $B$ are disjoint by item 4 of 4, the number of children $c$ of $s$ for which the set $(A\cup B)\cap V(T_{c})$ has odd size is $2$.
On the other hand, since $s\notin S^{\prime}$, 7(1) and (2) imply, using a similar argument to above, that there are no children $c$ of $s$ for which $(A\cup B)\cap V(T_{c})$ has odd size (in case (1)) or there is $1$ child $c$ of $s$ for which $(A\cup B)\cap V(T_{c})$ has odd size. This contradicts our conclusion from the previous paragraph that $s$ should have $2$ such children.
@@
Figure 4. Illustration of a bi-colouring $(A,B)$ identifying some set $S\subseteq V\setminus L$ in a binary tree $T$. Leaves from $A$ are blue diamonds, leaves from $B$ are orange filled circles, and inner nodes from $S$ are filled purple squares. Furthermore, the two paths connecting an inner node $s\in S$ to its AA- and $B$-representatives are coloured blue and orange, respectively. Nodes labeled $x_{1},x_{2},x_{3}$ constitute examples illustrating the different cases in 7.
@@
Proof 3.5.
If $T_{s}$ has height 0, then $s$ is a leaf and taking $t=s$ trivially gives the expected path. Otherwise, $s$ is an inner node and, because $X$ is thin, $s$ has at least one child $c_{s}$ not belonging to $X$. By induction, there is a path from some leaf $t\in V(T_{c_{s}})$ to $c_{s}$ avoiding $X$ and, since $s\notin X$, this path could be extended into a path from $t$ to $s$ avoiding $X$.
The following lemma allows to identify every thin set.
Lemma 10.
@@
Proof 3.6.
[…]
A family $F=(A_{1},B_{1}),\ldots,(A_{n},B_{n})$ of bi-colourings identifies a set $S\subseteq V\setminus L$, if there exists a partition $(S_{1},\ldots,S_{n})$ of $S$ such that, for each $i\in[n]$, $(A_{i},B_{i})$ identifies $S_{i}$. Whenever $S=V\setminus L$ we say that $F$ identifies $T$. Note that while $A_{i}$ and $B_{i}$ must be disjoint, for $i\not=j$ it is possible (and sometimes even necessary) that $A_{i}$ and $A_{j}$ or $A_{i}$ and $B_{j}$ are not disjoint (see e.g. Figure 5). A collection of subsets of $V\setminus L$ is thin if each of its subsets is thin. We now show that there exists a thin $4$-partition.
Lemma 11.
@@
Proof 3.8.
The result immediately follows from 10 and $1$.
Figure 5. Illustration of a partition of the inner nodes into 4 thin sets indicated by different shapes/fillings. The four colourings identifying each thin set are indicated in the table below the leaves where an orange slash indicates the leaf being in the corresponding set $A$ and a blue cross indicates the leaf being in $B$.
3.2. The transduction
The goal of this section is to prove 3, that is, to design a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction that produces the laminar tree induced by an input laminar set system. We fix a laminar set system $(U,\mathcal{S})$, represented by a $\{\mathsf{SET}\}$-structure $\mathbb{S}$. Before proving the theorem, we make the following basic observation. On $\mathbb{S}$, we can define two $\operatorname{MSO}$-formulas $\mathsf{desc}(X,Y)$ and $\mathsf{child}(X,Y)$ expressing that in the laminar tree induced by $\mathcal{S}$, $X$ and $Y$ are nodes and $X$ is a descendant or a child of $Y$, respectively:
@@ 3.2. The transduction
The key point of our construction consists in defining a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formula ${\mathsf{repr}}_{A,B}(a,X)$ which, assuming a bi-colouring $(A,B)$ of the universe modeled as disjoint unary relations and identifying a subset $S$ of inner nodes of $T$, is satisfied exactly when $X\in S$ and $a$ is its $A$-representative.
Lemma 13.
Let $(A,B)$ be a bi-colouring of $L(T)$ identifying a subset $S$ of inner nodes of $T$. There exists a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formula ${\mathsf{repr}}_{A,B}(a,X)$ that is satisfied exactly when $X$ is an inner node of $T$ that belongs to $S$ and is $A$-represented by $a$.
Proof 3.9.
@@
Now, we can easily define ${\mathsf{repr}}_{A,B}(a,X)$ based on 5:
${\mathsf{repr}}_{A,B}(a,X):=\>$ $\phi_{A,B}(X)\land a\in(X\cap A)\land\forall Z\subsetneq X\>\Big{(}a\in Z\rightarrow\neg\phi_{A,B}(Z)\Big{)}.$
@@
Proof 3.10 (Proof of 3).
The $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction is obtained by composing the following atomic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions. The transduction makes use of the formulas ${\mathsf{repr}}_{A,B}(a,X)$ given by 13.
  1. (1)
    Guess a family of four bi-colourings $(A_{i},B_{i})_{i\in[4]}$ identifying $T$ (which exists by 12).
  2. (2)
@@
  3. (3)
    Filter the universe keeping only the original elements as well as the $i$-th copy of each vertex $a$ for which there exists $X$ such that ${\mathsf{repr}}_{A_{i},B_{i}}(a,X)$.
  4. (4)
    Use an interpretation adding relation $\mathsf{ancestor}$ and keeping every other relation in $\Sigma$ unchanged outputting $\Sigma\cup\{\mathsf{ancestor}\}$-structures as follows. Define the relation $\mathsf{ancestor}(x,y)$ so that it is satisfied exactly when there exist $x^{\prime}$, $X$, $i$, and $Y$ such that, on the one hand $\mathsf{desc}(Y,X)$ and $\mathsf{copy}_{i}(x,x^{\prime})\land{\mathsf{repr}}_{A_{i},B_{i}}(x^{\prime},X)$, and, on the other hand, either $y$ is an original element and $Y=\{y\}$, or there exists $y^{\prime}$ and $j$ such that $\mathsf{copy}_{j}(y,y^{\prime})\land{\mathsf{repr}}_{A_{j},B_{j}}(y^{\prime},Y)$.
  5. (5)
    Use a filtering transduction keeping only those outputs $\mathbb{B}$ from the previous steps that satisfy that for every set $X$ in $\mathsf{SET}_{\mathbb{B}}$ there is an element $t$ in $U_{\mathbb{B}}$ such that the set of descendants of $t$ is equal to $X$ and for every element $t$ of $U_{\mathbb{B}}$ the set of descendants of $t$ is in $\mathsf{SET}_{\mathbb{B}}$ and hence verifying that $\mathsf{ancestor}_{\mathbb{B}}$ defines a laminar tree of $(U,\mathcal{S})$. Clearly, this can be expressed using a $\operatorname{MSO}$-sentence and hence outputs are of the form $\mathbb{A}\sqcup\mathbb{T}$ for some $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ representing the laminar tree $T$ of $(U,\mathcal{S})$.
4. Transducing modular decompositions
@@
4.1. Transducing weakly-partitive trees
@@
Figure 6. Sets to be included in a weakly-partitive set system (excluds the rightmost) or in a partitive set system (includes the rightmost) for two overlapping sets $X$ and $Y$.
A member of a set system $(U,\mathcal{S})$ is said to be strong if it does not overlap any other set from $\mathcal{S}$. The sub-family $\mathcal{S}_{!}$ of strong sets of $\mathcal{S}$ is thus laminar by definition. Hence, it induces a laminar tree $T$. By extension, we say that $T$ is induced by the set system $(U,\mathcal{S})$ (or simply by $\mathcal{S}$). The next result extends 3, by showing that $T$ can be $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduced from $\mathcal{S}$.
Lemma 14.
Let $\Sigma$ be an extended vocabulary, including a unary set predicate name $\mathsf{SET}$ and not including the binary relational symbol $\mathsf{ancestor}$. There exists a non-deterministic overlay $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each set system $(U,\mathcal{S})$ represented as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ and inducing the laminar tree $T$ with L(T)=U{L(T)=U}, and for each $\Sigma$-structure $\mathbb{A}$ with $\mathbb{S}\sqsubseteq\mathbb{A}$, $\tau(\mathbb{A})$ is non-empty and every output in $\tau(\mathbb{A})$ is equal to $\mathbb{A}\sqcup\mathbb{T}$ for some $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ representing $T$.
Proof 4.1.
@@
Hence, we can design an $\operatorname{MSO}$-interpretation that outputs the $\Sigma\cup\{\mathsf{SET!}\}$-structure corresponding to $\mathbb{A}$ equipped with the set unary predicate $\mathsf{SET!}$ that selects strong members of $\mathcal{S}$. Thus, up to renaming the predicates $\mathsf{SET!}$ and $\mathsf{SET}$, by 3, we can produce, through a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction, the $\Sigma\cup\{\mathsf{ancestor}\}$-structure $\mathbb{A}\sqcup\mathbb{T}$ where $\mathbb{T}$ is the tree-structure modeling the laminar tree $T$ induced by $\mathcal{S}_{!}$ with $L(T)=U$ (once the output obtained, the interpretation drops the set predicate $\mathsf{SET!}$ which is no longer needed).
@@
Clearly, the laminar tree $T$ of a weakly-partitive set system $(U,\mathcal{S})$ does not characterize $(U,\mathcal{S})$. However, as shown by the below theorem, a labeling of its inner nodes and a controlled partial ordering of its nodes are sufficient to characterize all the sets of $\mathcal{S}$. For $Z$ a set equipped with a partial order << and $X$ a subset of $Z$, we say that $X$ is a <<-interval whenever << defines a total order on $X$ and for every $a,b\in X$ and every $c\in Z$, $a<c<b$ implies $c\in X$.{thmC}[[CHM81, Rao06]] Let $\mathcal{S}$ be a weakly-partitive family, $\mathcal{S}_{!}$ be its subfamily of strong sets, and $T$ be the laminar tree it induces. There exists a total labeling function $\lambda$ from the set $V(T)\setminus L(T)$ of inner nodes of $T$ to the set $\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$, and, for each inner node $t\in\lambda^{-1}(\mathsf{linear})$, a linear ordering $_{t}$ of its children, such that every inner node having exactly two children is labeled by $\mathsf{degenerate}$ and the following two conditions are satisfied:
@@
  • for each $X\in\mathcal{S}\setminus\mathcal{S}_{!}$, there exists $t\in V(T)$ and a subset $\mathcal{C}$ of children of $t$ such that $X=\bigcup_{c\in\mathcal{C}}L(T_{c})$ and either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is a $_{t}-$interval, or $\lambda(t)=\mathsf{degenerate}$;
@@
  • conversely, for each inner node $t$ and each non-empty subset $\mathcal{C}$ of children of $t$, if either $\lambda(t)=\mathsf{linear}$ and $\mathcal{C}$ is a $_{t}-$interval, or $\lambda(t)=\mathsf{degenerate}$, then $\bigcup_{c\in\mathcal{C}}L(T_{c})\in\mathcal{S}$.
@@
Furthermore, $T$ and $\lambda$ are uniquely determined from $\mathcal{S}$, and, for each inner node $t$ of $T$ labeled by $\mathsf{linear}$, only two orders $_{t}$ are possible, one being the inverse of the other (indeed, inverting an order << does preserve the property of being a <<-interval). Hence, every weakly-partitive family $\mathcal{S}$ is characterized by a labeled and partially-ordered tree $(T,\lambda,<)$ where $T$ is the laminar tree induced by the subfamily $\mathcal{S}_{!}$ of strong sets of $\mathcal{S}$, $\lambda:V(T)\setminus L(T)\to\{\mathsf{degenerate},\mathsf{prime},\mathsf{linear}\}$ is the labeling function, and << is the partial order $\bigcup_{t\in\lambda^{-1}(\mathsf{linear})}<_{t}$ over $V(T)$. As, up-to inverting some of the $_{t}$ orders, $(T,\lambda,<)$ is unique, we abusively call it the weakly-partitive tree induced by $\mathcal{S}$. Conversely, a weakly-partitive tree characterizes the unique weakly-partitive set system which induced it.
@@
We naturally model a weakly-partitive tree $(T,\lambda,<)$ of a weakly-partitive set system $(U,\mathcal{S})$ by the $\{\mathsf{ancestor},\mathsf{degenerate},\mathsf{betweeness}\}$-structure $\mathbb{T}$ of universe $U_{\mathbb{T}}=V(T)$ such that $\langle V(T),\mathsf{ancestor}_{\mathbb{T}}\rangle\sqsubseteq\mathbb{T}$ models $T$ with L(T)=U{L(T)=U}, $\mathsf{degenerate}_{\mathbb{T}}$ is a unary relation which selects the inner nodes of $T$ of label $\mathsf{degenerate}$, i.e., $\mathsf{degenerate}_{\mathbb{T}}=\lambda^{-1}(\mathsf{degenerate})$}, and $\mathsf{betweeness}_{\mathbb{T}}$ is a ternary relation selecting triples $(x,y,z)$ satisfying $x<y<z$ or $z<y<x$. (Although it is possible, through a non-deterministic $\operatorname{MSO}$-transduction, to define << from $\mathsf{betweeness}_{\mathbb{T}}$, the use of $\mathsf{betweeness}_{\mathbb{T}}$ rather than $_{\mathbb{T}}$ ensures uniqueness of the output weakly-partitive tree.) The inner nodes of $T$ that are labeled by $\mathsf{linear}$ can be recovered through an $\operatorname{MSO}$-formula as those inner nodes whose children are related by $\mathsf{betweeness}$. The inner nodes labeled by $\mathsf{prime}$ can be recovered through an $\operatorname{MSO}$-formula as those inner nodes that are labeled neither by $\mathsf{degenerate}$ nor by $\mathsf{linear}$. Using 4.1, it is routine to design an $\operatorname{MSO}$-transduction which takes as input a weakly-partitive tree and outputs the weakly-partitive set system which induced it. The inverse $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction is the purpose of the next result.
Theorem 15.
@@
Proof 4.2.
By 14, we have a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which on inputs $\mathbb{S}$ outputs the $\{\mathsf{SET},\mathsf{ancestor}\}$-structure $\mathbb{S}\sqcup\mathbb{T^{\prime}}$, where $\mathbb{T^{\prime}}$ is the $\{\mathsf{ancestor}\}$-structure modeling $T$, with $L(T)=U=U_{\mathbb{S}}$. It thus remains to define a second $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction which outputs $\mathbb{T}$ from $\mathbb{S}\sqcup\mathbb{T^{\prime}}$. Actually, an $\operatorname{MSO}$-interpretation is enough. Indeed, we can define an $\operatorname{MSO}$-formula $\phi_{\mathsf{degenerate}}$ which selects the inner nodes of $T$ that are labeled $\mathsf{degenerate}$ by $\lambda$, and an $\operatorname{MSO}$-formula $\phi_{\mathsf{betweeness}}$ which selects triples $(x,y,z)$ of elements such that $x<y<z$ or $z<y<x$. In order to define both formulas, we use a functional symbol $\mathsf{leafset}(t)$ which is interpreted as the set $L(T_{t})$, and is clearly $\operatorname{MSO}$-definable on $\{\mathsf{ancestor}\}$-structures such as $\mathbb{T}$. Indeed, an element $z$ belongs to $\mathsf{leafset}(t)$ if and only if it is a leaf and has $t$ as ancestor.
According to 4.1, an inner node $t$ is labeled $\mathsf{degenerate}$ if and only if, for every two children $s_{1},s_{2}$ of $t$ the set $L(T_{s_{1}})\cup L(T_{s_{2}})$ is in $\mathcal{S}$. We can define this in $\operatorname{MSO}$ using the following formula:
@@
Now, according to 4.1, we have $x<y<z$ or $z<y<x$ exactly when $x$, $y$, and $z$ are three distinct children of some node $t$ and the three following properties are satisfied:
  1. (1)